OpenStudy (anonymous):

What velocity does a freely falling object reach after 4.0 s if it starts from rest?

OpenStudy (zephyr141):

\[v=v_{0}+at\]use this. and remember that "starts from rest" means that's initial velocity is going to be 0 m/s and that the acceleration is the acceleration of gravity on earth so -9.8 m/s^2 and just plug and chug.

OpenStudy (anonymous):

so -9.8 m/s multiplied by time right?

OpenStudy (zephyr141):

yeah that will give you your answer.

OpenStudy (anonymous):

thank you!!!!

OpenStudy (zephyr141):

no problem! keep working hard!

OpenStudy (anonymous):

but wait will it be negative or positive? does it depend now? Can it be negative?

OpenStudy (zephyr141):

well it's going to be the positive number if you're just looking at velocity. but the negative sign in this case just gives us the direction that you're accelerating in. which is down.

OpenStudy (zephyr141):

soon you'll be looking at force vectors and such which have direction so the sign of most of the answers that you'll be getting are just meant as orientation to know which direction that force is traveling in.

OpenStudy (anonymous):

ok thank you!!! *your a better teacher than my physics one....*

OpenStudy (zephyr141):

lol always ask your teacher for help. that will always help you. if you have a question don't be afraid to ask. teachers are there to help. also i struggle at physics so i know how it is for others and can convey the content sometimes. but really, the teacher will know more and would more than likely be glad to help you. also just keep chuggin along and you'll get there.

OpenStudy (anonymous):

\[v =u +at ,a =+ g, u =0, t =4, v =0+4g =40m/s\]